# The Mixed Birth-death/death-Birth Moran Process

### v2

@@ 1 Introduction
Section˜2 formalizes the mixed Moran process and describes some technical tools. Section˜3 analyzes the unbiased mixed Moran process where $\lambda=1/2$. Section˜4 examines almost-regular and random graphs. Section˜5 studies graphs with two values of degrees. Explicit formulas for the stars and the cycles appear in Section˜6. Section˜7 contains some concluding remarks. Omitted proof details can be found in the Appendices.
@@ 2.1 Models and Notation
Before delving into specific regimes, we provide two general characteristics for the fixation probabilities in the mixed Moran process. These results were previously known for the pure Bd and dB update rules, and we generalize them to the mixed process. The proofs appear in Section˜A.1.
@@
We believe that one could prove a stronger statement that the fixation probability is monotone in $r$ and the current set $S$. While seemingly intuitive, there is actually some subtlety in proving it rigorously (cf. Remark˜A.1 for more discussions on this). We leave proving the property for future work.
@@ 2.2 Technical tools
We use a potential-function (drift) method. Fix a potential function $\phi$ on subsets $S\subseteq V$. Our goal is to bound the expected one-step change of $\phi$. These theorems are relatively standard, so we defer the proofs to Section˜A.2.
@@ 3 The case $\lambda=1/2$
$\psi_{u,v}(S)$ $=\frac{1}{2}\left[\frac{1}{n}\cdot\frac{1}{\deg(u)}-\frac{1}{n}\cdot\frac{1}{\deg(v)}\right]+\frac{1}{2}\left[\frac{1}{n}\cdot\frac{1}{\deg(v)}-\frac{1}{n}\cdot\frac{1}{\deg(u)}\right]=0.$
@@
The claim follows from Theorem˜2.6 and 2.3. ∎
@@
Because $\phi(S)=0$ at $S=\emptyset$ and $\phi(S)=n$ at $S=V$, applying Theorem˜2.5 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ gives
@@
Recall that the fixation probability with initial mutants $S_{0}$ is $|S_{0}|/n$ when $r=1$. Combined with Theorem˜2.4, we yield the following corollary.
@@
Applying Theorem˜2.7 together with Theorem˜3.3 and Corollary˜3.4 gives an FPRAS for Mixed Moran fixation on any graph when $\lambda=1/2$ and $r\geqslant 1$.
@@ 4 Almost Regular Graphs and Random Graphs
We first show that for a regular graph $G$, the fixation probability starting with an initial mutant set $S_{0}$ when $r=1$ is $|S_{0}|/n$, regardless of $\lambda$ and the position of the mutants. This could be compared with Theorem˜3.1 where the same property holds for $\lambda=1/2$ and all graphs: both cases demonstrate this property because the process balances the change in different directions, resulting in the noted invariance. Similar to Theorem˜3.3, we also have the absorption time is $O_{r}(n^{4})$. Since the proofs are very similar, we defer them to Appendix˜B.
@@
It is somewhat a strong requirement to require regularity, and we may expect that if the graph is “almost regular", things may still behave nicely.
@@
Recall from Section˜3 that $\psi^{1/2}_{u,v}(S)\geqslant(r-1)/(rn^{3})$. We now bound $\psi^{Bd}_{u,v}$.
@@
We may assume the graph is not regular, since otherwise the result is already implied by Theorem˜4.1. Thus, $\alpha\geqslant(n-1)/(n-2)$. For $n\geqslant 3$, since $r\geqslant\alpha^{2}$ and $\frac{n-1}{n-2}\geqslant\frac{n}{n-1}$,
@@
Applying Theorem˜2.5 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$.
For the fixation probability, applying Theorem˜2.6 with $k_{0}=1$ and $k_{1}=n$ gives $\mathrm{fp}(S_{0})\geqslant\frac{1}{n}$}.
@@
We call a configuration bad if every mutant has only resident neighbors and every resident has only mutant neighbors, i.e., for all $u\in S$, $N(u)\subseteq V\setminus S$ and for all $v\notin S$, $N(v)\subseteq S$. In a bad state, the mutant has no selection advantage (in Equation˜1, we see the $r$ in the numerator and denominator cancel out). Accordingly, we have no strictly positive lower bounds on the drift.  [^5] It may seem counterintuitive that a configuration is bad when all neighbors are of different types: in this case, the network is guaranteed to evolve instead of self-looping. However, while the state changes, we cannot ensure that it (in expectation) moves in the desired direction (i.e., increasing the potential function). Conversely, when some neighbors are of the same type, the system is more likely to remain in the same configuration, but whenever it changes, it tends to move in a favorable direction.
@@
Applying Theorem˜2.5 with $k_{1}=n^{2}$, and $k_{2}=(r-1)/(2rn^{2})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$. For the fixation probability, applying Theorem˜2.6 with $k_{0}\geqslant 1$ and $k_{1}=n^{2}$ gives $\mathrm{fp}(S_{0})\geqslant n^{-2}$}. ∎
@@
Combining Theorem˜2.7, 4.3, and 4.5 yields the FPRAS for random graphs.
@@ 5 Bidegreed Graphs
In this section and the following section, we consider some interesting special cases of graphs. Here we consider graphs where the degrees all take on one of two possible values, also called bidegreed graphs (e.g., [myrvold1987bidegreed, belardo2009bidegreed]). This class of graphs is interesting for two reasons. First, in some sense, it is another way of examining graphs that are nearly regular, albeit in a different way than our $\alpha$-almost regular graphs. Additionally, several natural graph types, including stars and bipartite biregular graphs, are bidegreed; see Figure˜1.
@@
We note that a $(d_{1},d_{2})$-bidegreed graph is also $(d_{2}/d_{1})$-almost regular. Thus, the results from Section˜4 apply. However, with the extra guarantee that all vertices have degrees that take on only two possible values, we achieve stronger results. We first show a closed-form expression for the fixation probability under $r=1$.
@@
We show the result for singleton sets $S=\{u\}$; the general theorem follows from Theorem˜2.3. We use the potential function $\phi(u)=f(\deg_{u})$. Now fix $S$ and consider $(u,v)\in\mathrm{bdry}(S)$. We have
@@
and $\phi(S_{i})$ is a martingale. Applying Theorem˜2.6 with $k_{0}=f(\deg_{u})$ and $k_{1}=\sum_{v}f(\deg_{v})$ gives the theorem. ∎
Next, we establish a bound for the absorption time. We could apply Theorem˜4.3 when $r\geqslant\alpha^{2}=(d_{2}/d_{1})^{2}$. However, for bidegreed graphs, we can remove this assumption and bound the absorption time for any $r$. We remark, however, that the exponent in our bound is probably not tight; we leave proving better bounds for future work.
@@
We do a case analysis based on whether $r=1$. We employ the potential $\phi(u)=f(\deg_{u})$ with $f$ the same as Theorem˜5.2.
@@
Recall from Theorem˜5.2 that the neutral terms cancel each other:
@@
Recall from the proof of Theorem˜4.3 that we say a configuration $S$ is bad if $N(u)\subseteq V\setminus S$ for all $u\in S$ and $N(v)\subseteq S$ for all $v\notin S$. In that case, $|N(v)\setminus S|=|N(u)\cap S|=0$ and we have no strictly positive lower bound for dB. Consequently, we only have $\psi^{\lambda}_{u,v}(S)\geqslant 0$, but we can use that $S$ will transit into a non-bad configuration in one step.
@@
Using the same idea as in Theorem˜4.3 of collapsing a step from a bad configuration into the next step, and applying Theorem˜2.5 with $k_{1}=n\alpha$ and $k_{2}=\frac{r-1}{rn^{3}\alpha}$, the absorption time is bounded by $O_{r}(n^{4}\alpha^{2})$.
@@
As shown in Theorem˜5.2, $\mathbb{E}\!\left[\phi(S_{i+1})-\phi(S_{i})\mid S_{i}\right]=0$. Let
@@
so $(Z_{i})$ is a supermartingale with bounded one–step differences. Similar to the proof of Theorem˜3.3, it is easy to verify that $\tau$ has finite expectation. Thus, applying optional stopping at the absorption time $\tau$,
@@ 6.1 Cycles
Note that cycles are $2$-regular graphs, thus all results from Section˜4 apply. On top of that, we can solve for the exact fixation probability for any $r$ and the graph starts with one mutant vertex. In this setting, we only need to control the current number of mutants, which are contiguous on the cycle. We compute the probability of the mutant size increasing $p_{k}^{\uparrow}$ or decreasing $p_{k}^{\downarrow}$ when there are $k$ mutants.
Figure 2: Fixation probabilities for a cycle on $n=100$ vertices starting from a single mutant. Fixation probabilities are computed using Equation˜2. The two parameters are $r$ and $\lambda$, with $\lambda$ encoded by color. When $r<1$, the fixation probabilities are very small (but not actually 0).
@@ 6.2 Stars
Figure 3: Fixation probabilities for a star on $10$ vertices starting from a single initial mutant, computed using the results from Section˜6.2. The two parameters are $r$ and $\lambda$, with $\lambda$ encoded by color. When the initial mutant is at the center, larger $\lambda$ (more Bd) consistently yields a smaller fixation probability. When starting from a leaf, there appears to be a phase transition around $r=1$ about whether greater Bd or dB gives a higher fixation probability.
@@ A.1 Basic Properties of mixed Moran Process
We use a coupling argument. Let $(Z_{t})_{t\geqslant 0}$ denote the neutral process with fitness $1$ and $(X_{t})_{t\geqslant 0}$ denote the advantageous process with fitness $r>1$, both starting from the same set $Z_{0}=X_{0}=S$. We construct a coupling such that $Z_{t}\subseteq X_{t}$ for all steps $t$. Consider the transition from time step $t$ to $t+1$ and any vertex $v\in V$.
  • We bound the probability that a mutant is placed at $v$. Let $p^{Z}_{t+1}(v)$ denote the probability that a mutant is placed at $v$ in $Z_{t+1}$, conditioned on $Z_{t}$, and $p^{X}_{t+1}(v)$ denote the same quantity for process $(X_{t})$. We have
    $p^{Z}_{t+1}(v)$ $=\lambda\sum_{u\in N(v)\cap Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{u}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{|N(v)\cap Z_{t}|}{\deg_{v}}$
    $p^{X}_{t+1}(v)$ $=\lambda\sum_{u\in N(v)\cap X_{t}}\frac{r}{w(X_{t})}\cdot\frac{1}{\deg_{u}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{r|N(v)\cap X_{t}|}{w_{v}(X_{t})}$
    Since $r\geqslant 1$ and $Z_{t}\subseteq X_{t}$, we have
    $\sum_{u\in N(v)\cap Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{u}}\leqslant$ $\sum_{u\in N(v)\cap X_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{u}}\leqslant\sum_{u\in N(v)\cap X_{t}}\frac{r}{w(X_{t})}\cdot\frac{1}{\deg_{u}},$
    $\frac{1}{n}\cdot\frac{|N(v)\cap Z_{t}|}{\deg_{v}}\leqslant$ $\frac{1}{n}\cdot\frac{|N(v)\cap X_{t}|}{\deg_{v}}\leqslant\frac{1}{n}\cdot\frac{r|N(v)\cap X_{t}|}{w_{v}(X_{t})},$
    and therefore $p^{Z}_{t+1}(v)\leqslant p^{X}_{t+1}(v)$.
  • We bound the probability that a resident is placed at $v$. Let $q^{Z}_{t+1}(v)$ denote the probability that a resident is placed at $v$ in $Z_{t+1}$, conditioned on $Z_{t}$, and $q^{X}_{t+1}(v)$ denote the same quantity for process $(X_{t})$. Then,
    $q^{Z}_{t+1}(v)$ $=\lambda\sum_{v\in N(v)\setminus Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{v}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{|N(v)\setminus Z_{t}|}{\deg_{u}},$
    $q^{X}_{t+1}(v)$ $=\lambda\sum_{v\in N(v)\setminus X_{t}}\frac{1}{w(X_{t})}\cdot\frac{1}{\deg_{v}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{|N(v)\setminus X_{t}|}{w_{u}(X_{t})}$
    Since $r\geqslant 1$, we have $w(X_{t})\geqslant n$ and $w_{u}(X_{t})\geqslant\deg_{u}$. Furthermore, $N(v)\setminus Z_{t}\supseteq N(v)\setminus X_{t}$. Thus,
    $\sum_{v\in N(v)\setminus Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{v}}$ $\geqslant\sum_{v\in N(v)\setminus X_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{v}}\geqslant\sum_{v\in N(v)\setminus X_{t}}\frac{1}{w(X_{t})}\cdot\frac{1}{\deg_{v}},$
    $\frac{1}{n}\cdot\frac{|N(v)\setminus Z_{t}|}{\deg_{u}}$ $\geqslant\frac{1}{n}\cdot\frac{|N(v)\setminus X_{t}|}{\deg_{u}}\geqslant\frac{1}{n}\cdot\frac{|N(v)\setminus X_{t}|}{w_{u}(X_{t})},$
    and therefore $q^{Z}_{t+1}(v)\geqslant q^{X}_{t+1}(v)$.
Thus, we can couple the transitions such that if a mutant is placed at vertex $v$ in $(Z_{t})$, a mutant is also placed at $v$ in $(X_{t})$ [^6] Note that it could be the case that $Z_{t}$ or $X_{t}$ already contains $v$. That is fine and does not affect our coupling.; if a resident is placed at vertex $v$ in $(X_{t})$, a resident is also placed at $v$ in $(Z_{t})$. Consequently, $Z_{t}\subseteq X_{t}$ holds for all $t$. We conclude that $\mathrm{fp}^{\lambda,1}_{G}(S)\leqslant\mathrm{fp}^{\lambda,r}_{G}(S)$. ∎
@@
However, for the mixed process, this solution faces another challenge: In the continuous embedding for Bd, each vertex has a rate of $f(u)$ for Birth, whereas dB has a fixed rate of $1$ for death. This is problematic since this render will break the condition that the ratio between the rates of Bd and dB events remains $\lambda/(1-\lambda)$; otherwise, the discretization of the continuous-time chain is not equivalent to the discrete Moran process anymore. A natural attempt is then to scale the dB rate by the total fitness to fix this ratio, but this brings back the original problem: having more mutants increases the total fitness and thus the rate of death events, which again breaks the natural coupling. Regardless, we believe that monotonicity in $r$ still holds, although establishing it likely requires a more careful analysis.
A.2 Martingales and Drifts

### v3

@@ 1 Introduction
Section 2 formalizes the mixed Moran process and describes some technical tools. Section 3 analyzes the unbiased mixed Moran process where $\lambda=1/2$. Section 4 examines almost-regular and random graphs. Section 5 studies graphs with two values of degrees. Explicit formulas for the stars and the cycles appear in Section 6. Section 7 contains some concluding remarks. Omitted proof details can be found in the Appendices.
@@ 2.1 Models and Notation
Before delving into specific regimes, we provide two general characteristics for the fixation probabilities in the mixed Moran process. These results were previously known for the pure Bd and dB update rules, and we generalize them to the mixed process. The proofs appear in Section A.1.
@@
We believe that one could prove a stronger statement that the fixation probability is monotone in $r$ and the current set $S$. While seemingly intuitive, there is actually some subtlety in proving it rigorously (cf. Remark A.1 for more discussions on this). We leave proving the property for future work.
@@ 2.2 Technical tools
We use a potential-function (drift) method. Fix a potential function $\phi$ on subsets $S\subseteq V$. Our goal is to bound the expected one-step change of $\phi$. These theorems are relatively standard, so we defer the proofs to Section A.2.
@@ 3 The case $\lambda=1/2$
$\psi_{u,v}(S)$ $=\frac{1}{2}\left[\frac{1}{n}\cdot\frac{1}{\deg_{u}}-\frac{1}{n}\cdot\frac{1}{\deg_{v}}\right]+\frac{1}{2}\left[\frac{1}{n}\cdot\frac{1}{\deg_{v}}-\frac{1}{n}\cdot\frac{1}{\deg_{u}}\right]=0.$
@@
The claim follows from Theorem 2.6 and 2.3. ∎
@@
Because $\phi(S)=0$ at $S=\emptyset$ and $\phi(S)=n$ at $S=V$, applying Theorem 2.5 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ gives
@@
Recall that the fixation probability with initial mutants $S_{0}$ is $|S_{0}|/n$ when $r=1$. Combined with Theorem 2.4, we yield the following corollary.
@@
Applying Theorem 2.7 together with Theorem 3.3 and Corollary 3.4 gives an FPRAS for Mixed Moran fixation on any graph when $\lambda=1/2$ and $r\geqslant 1$.
@@ 4 Almost Regular Graphs and Random Graphs
We first show that for a regular graph $G$, the fixation probability starting with an initial mutant set $S_{0}$ when $r=1$ is $|S_{0}|/n$, regardless of $\lambda$ and the position of the mutants. This could be compared with Theorem 3.1 where the same property holds for $\lambda=1/2$ and all graphs: both cases demonstrate this property because the process balances the change in different directions, resulting in the noted invariance. Similar to Theorem 3.3, we also have the absorption time is $O_{r}(n^{4})$. Since the proofs are very similar, we defer them to Appendix B.
@@
It is somewhat a strong requirement to require regularity, and we may expect that if the graph is “almost regular”, things may still behave nicely.
@@
Recall from Section 3 that $\psi^{1/2}_{u,v}(S)\geqslant(r-1)/(rn^{3})$. We now bound $\psi^{Bd}_{u,v}$.
@@
We may assume the graph is not regular, since otherwise the result is already implied by Theorem 4.1. Thus, $\alpha\geqslant(n-1)/(n-2)$. For $n\geqslant 3$, since $r\geqslant\alpha^{2}$ and $\frac{n-1}{n-2}\geqslant\frac{n}{n-1}$,
@@
Applying Theorem 2.5 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$.
For the fixation probability, applying Theorem 2.6 with $k_{0}=1$ and $k_{1}=n$ gives $\mathrm{fp}(S_{0})\geqslant\frac{1}{n}$}.
@@
We call a configuration bad if every mutant has only resident neighbors and every resident has only mutant neighbors, i.e., for all $u\in S$, $N(u)\subseteq V\setminus S$ and for all $v\notin S$, $N(v)\subseteq S$. In a bad state, the mutant has no selection advantage (in Equation 1, we see the $r$ in the numerator and denominator cancel out). Accordingly, we have no strictly positive lower bounds on the drift.  [^5] It may seem counterintuitive that a configuration is bad when all neighbors are of different types: in this case, the network is guaranteed to evolve instead of self-looping. However, while the state changes, we cannot ensure that it (in expectation) moves in the desired direction (i.e., increasing the potential function). Conversely, when some neighbors are of the same type, the system is more likely to remain in the same configuration, but whenever it changes, it tends to move in a favorable direction.
@@
Applying Theorem 2.5 with $k_{1}=n^{2}$, and $k_{2}=(r-1)/(2rn^{2})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$. For the fixation probability, applying Theorem 2.6 with $k_{0}\geqslant 1$ and $k_{1}=n^{2}$ gives $\mathrm{fp}(S_{0})\geqslant n^{-2}$}. ∎
@@
Combining Theorem 2.7, 4.3, and 4.5 yields the FPRAS for random graphs.
@@ 5 Bidegreed Graphs
In this section and the following section, we consider some interesting special cases of graphs. Here we consider graphs where the degrees all take on one of two possible values, also called bidegreed graphs (e.g., [myrvold1987bidegreed, belardo2009bidegreed]). This class of graphs is interesting for two reasons. First, in some sense, it is another way of examining graphs that are nearly regular, albeit in a different way than our $\alpha$-almost regular graphs. Additionally, several natural graph types, including stars and bipartite biregular graphs, are bidegreed; see Figure 1.
@@
We note that a $(d_{1},d_{2})$-bidegreed graph is also $(d_{2}/d_{1})$-almost regular. Thus, the results from Section 4 apply. However, with the extra guarantee that all vertices have degrees that take on only two possible values, we achieve stronger results. We first show a closed-form expression for the fixation probability under $r=1$.
@@
We show the result for singleton sets $S=\{u\}$; the general theorem follows from Theorem 2.3. We use the potential function $\phi(u)=f(\deg_{u})$. Now fix $S$ and consider $(u,v)\in\mathrm{bdry}(S)$. We have
@@
and $\phi(S_{i})$ is a martingale. Applying Theorem 2.6 with $k_{0}=f(\deg_{u})$ and $k_{1}=\sum_{v}f(\deg_{v})$ gives the theorem. ∎
Next, we establish a bound for the absorption time. We could apply Theorem 4.3 when $r\geqslant\alpha^{2}=(d_{2}/d_{1})^{2}$. However, for bidegreed graphs, we can remove this assumption and bound the absorption time for any $r$. We remark, however, that the exponent in our bound is probably not tight; we leave proving better bounds for future work.
@@
We do a case analysis based on whether $r=1$. We employ the potential $\phi(u)=f(\deg_{u})$ with $f$ the same as Theorem 5.2.
@@
Recall from Theorem 5.2 that the neutral terms cancel each other:
@@
Recall from the proof of Theorem 4.3 that we say a configuration $S$ is bad if $N(u)\subseteq V\setminus S$ for all $u\in S$ and $N(v)\subseteq S$ for all $v\notin S$. In that case, $|N(v)\setminus S|=|N(u)\cap S|=0$ and we have no strictly positive lower bound for dB. Consequently, we only have $\psi^{\lambda}_{u,v}(S)\geqslant 0$, but we can use that $S$ will transit into a non-bad configuration in one step.
@@
Using the same idea as in Theorem 4.3 of collapsing a step from a bad configuration into the next step, and applying Theorem 2.5 with $k_{1}=n\alpha$ and $k_{2}=\frac{r-1}{rn^{3}\alpha}$, the absorption time is bounded by $O_{r}(n^{4}\alpha^{2})$.
@@
As shown in Theorem 5.2, $\mathbb{E}\!\left[\phi(S_{i+1})-\phi(S_{i})\mid S_{i}\right]=0$. Let
@@
so $(Z_{i})$ is a supermartingale with bounded one–step differences. Similar to the proof of Theorem 3.3, it is easy to verify that $\tau$ has finite expectation. Thus, applying optional stopping at the absorption time $\tau$,
@@ 6.1 Cycles
Note that cycles are $2$-regular graphs, thus all results from Section 4 apply. On top of that, we can solve for the exact fixation probability for any $r$ and the graph starts with one mutant vertex. In this setting, we only need to control the current number of mutants, which are contiguous on the cycle. We compute the probability of the mutant size increasing $p_{k}^{\uparrow}$ or decreasing $p_{k}^{\downarrow}$ when there are $k$ mutants.
Figure 2: Fixation probabilities for a cycle on $n=100$ vertices starting from a single mutant. Fixation probabilities are computed using Equation 2. The two parameters are $r$ and $\lambda$, with $\lambda$ encoded by color. When $r<1$, the fixation probabilities are very small (but not actually 0).
@@ 6.2 Stars
Figure 3: Fixation probabilities for a star on $10$ vertices starting from a single initial mutant, computed using the results from Section 6.2. The two parameters are $r$ and $\lambda$, with $\lambda$ encoded by color. When the initial mutant is at the center, larger $\lambda$ (more Bd) consistently yields a smaller fixation probability. When starting from a leaf, there appears to be a phase transition around $r=1$ about whether greater Bd or dB gives a higher fixation probability.
@@ A.1 Basic Properties of mixed Moran Process
We use a coupling argument. Let $(Z_{t})_{t\geqslant 0}$ denote the neutral process with fitness $1$ and $(X_{t})_{t\geqslant 0}$ denote the advantageous process with fitness $r>1$, both starting from the same set $Z_{0}=X_{0}=S$. We construct a coupling such that $Z_{t}\subseteq X_{t}$ for all steps $t$. Consider the transition at time $t$. For $u\in V,v\in N(u)$, we denote by $q_{t}^{Z}(u\to v)$ the probability that vertex $u$ is selected for birth and replaces neighbour $v$ in $Z_{t}$, and $q_{t}^{Z}(u\leftarrow v)$ the probability that vertex $u$ is selected for death and is replaced by neighbour $v$ in $Z_{t}$. Similarly, $q_{t}^{X}$ denotes the same quantity for the chain $(X_{t})$.
  • If $Z_{t}$ performs a Bd update with a mutant $u$ replacing a neighbour $v$, $X_{t}$ also performs a Bd update with mutant $u$ replacing $v$. This is possible since
    $q_{t}^{Z}(u\to v)=\frac{\lambda}{n\deg_{u}}\leqslant\frac{\lambda}{\deg_{u}}\frac{r}{w(X_{t})}=q_{t}^{X}(u\to v).$
  • If $X_{t}$ performs a Bd update with resident $v$ replacing a neighbour $u$, $Z_{t}$ also performs a Bd update with resident $v$ replacing $u$. This is possible since
    $q_{t}^{X}(v\to u)=\frac{\lambda}{\deg_{v}}\frac{1}{w(X_{t})}\leqslant\frac{\lambda}{n\deg_{v}}=q_{t}^{Z}(v\to u).$
  • If $Z_{t}$ performs a dB update with a vertex $v$ replaced by a mutant neighbour $u$, $X_{t}$ also does that. This is possible since
    $q_{t}^{Z}(v\leftarrow u)=\frac{1-\lambda}{n\deg_{v}}\leqslant\frac{1-\lambda}{n}\frac{r}{w_{v}(X_{t})}=q_{t}^{X}(v\leftarrow u).$
  • If $X_{t}$ performs a dB update with a vertex $u$ replaced by a resident neighrbour $v$, $Z_{t}$ also does that. This is possible since
    $q_{t}^{X}(u\leftarrow v)=\frac{1-\lambda}{n}\frac{1}{w_{v}(X_{t})}\leqslant\frac{1-\lambda}{n\deg_{v}}=q_{t}^{Z}(u\leftarrow v).$
  • All remaining probabilities place a mutant in $X_{t}$ and a resident in $Z_{t}$, so we can couple them arbitrarily.
Following this coupling, $Z_{t}\subseteq X_{t}$ holds for all $t$. We conclude that $\mathrm{fp}^{\lambda,1}_{G}(S)\leqslant\mathrm{fp}^{\lambda,r}_{G}(S)$. ∎
@@
However, for the mixed process, this solution faces another challenge: In the continuous embedding for Bd, each vertex has a rate of $f(u)$ for Birth, whereas dB has a fixed rate of $1$ for death. This is problematic since it breaks the condition that the ratio between the rates of Bd and dB events remains $\lambda/(1-\lambda)$ and hence the discretization of the continuous-time chain is not equivalent to the discrete Moran process anymore. A natural attempt is then to scale the dB rate by the total fitness to fix this ratio, but this brings back the original problem: having more mutants increases the total fitness and thus the rate of death events, which again breaks the natural coupling. Regardless, we believe that monotonicity in $r$ still holds, although establishing it likely requires a more careful analysis.
A.2 Martingales, Drifts, and Approximation Schemes